OpenStudy (anonymous):

Find the value of 7 times x squared all over the square root of x to the fifth power.

OpenStudy (anonymous):

\[\frac{ 7x^2 }{ \sqrt{x^5} }\]

OpenStudy (anonymous):

does it look like that?

OpenStudy (anonymous):

yes!

OpenStudy (anonymous):

first we need to write \(\sqrt{x^5}\) in simplest radical form since 2 goes in to 5 twice, with a remainder of 1, we get \[\sqrt{x^5}=x^2\sqrt{x}\]

OpenStudy (anonymous):

Hey satellite when your done here can you help me with my problem. No one has answered it in 14 hours. Thank you

OpenStudy (anonymous):

@iamlegend i can try, lets finish this one

OpenStudy (anonymous):

we get \[\frac{ 7x^2 }{ \sqrt{x^5} }=\frac{7x^2}{x^2\sqrt{x}}\] and now we can cancel top and bottom to get \[\frac{7x^2}{x^2\sqrt{x}}=\frac{7}{\sqrt{x}}\]

OpenStudy (anonymous):

if you have to write this in simplest radical form, with no radicals in the denominator, it is \[\frac{7}{\sqrt{x}}\times \frac{\sqrt{x}}{\sqrt{x}}=\frac{7\sqrt{x}}{x}\]

OpenStudy (anonymous):

That was very understandable. Thank you so much!

OpenStudy (anonymous):

yw